Suppose the demand for a product is given by Qc=100 - 2Pt+Pc where Qc is the quantity demand for coffee,Pc is the price of coffee,Pt is the price of the substitute good tea.The price of the substitute good is ₹400/kg.Suppos Pc=₹650/kg,what is the price elasticity of demand?


1
Expert's answer
2021-02-25T18:24:53-0500

Qc= 100-2Pt+Pc

Therefore,"\frac{dQ_c}{dP_c} = 1"

Qo= 100-2(400)+650= -50

"P_t= 650"

"PED=-\frac{dQ}{dP} \times \frac{P_t}{Q_o}\\

\implies -1 \times \frac{650}{-50}\\

\implies 13.\\
\text{The demand is elastic.}"